MICROWAVE SPECTRA AND MOLECULAR STRUCTURES OF 2-CHLORO-1,1-DIFLUOROETHYLENE AND ITS COMPLEX WITH ARGON

\underline{JOSEPH P. MESSINGER}, GREGORY S. KNOWLTON, KATHRYN M. SUNDHEIM, HE-LEN O. LEUNG, AND MARK D. MARSHALL, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

Chirped-pulse and Balle-Flygare spectrometers are used to obtain Fourier transform microwave spectra of 2-chloro-1,1-difluoroethylene and its complex with argon from 5.5 to 21.0 GHz, allowing for the geometries of both species to be determined. A total of six isotopologues are observed each for the monomer and dimer, including the most abundant species, the singly-substituted ^{37}Cl and two singly-substituted ^{13}C isotopologues in natural abundance, and deuterated versions of both the ^{35}Cl and ^{37}Cl species using an isotopically enriched sample. Similar to the previously studied argonhaloethylene complexes, the argon shows a preference for close contact with heavier atoms. Tunneling of the argon between two equivalent non-planar structures, similar to that in argon-cis-1,2-difluoroethylene, is not observed in this complex.